# Targeted Deep Learning System Boundary Testing

### v2

@@ 1. Introduction
The increasing dependence on Deep Learning (DL) systems for both everyday tasks and critical sectors (Schütze et al., 2008) makes rigorous testing for these systems a relevant topic (Riccio et al., 2020; Zhang et al., 2020b). The concept of fault in DL systems is more complex than in traditional software (Riccio et al., 2020). Even if the code that builds the DL network is bug-free, the trained DL model may still deviate from the expected behavior due to faults introduced during the training phase, such as the misconfiguration of learning parameters or the use of an unbalanced or non-representative training set (Humbatova et al., 2020). In data-intensive software systems, such as DL systems, faults often stem from the large, high-dimensional input space, which requires the generation of test data that accurately captures the complexity and diversity of the validity domain, i.e., the portion of the input space for which the system is expected to operate (Riccio et al., 2020).
@@
Test generation techniques have been developed to induce misbehaviors in DL systems (Riccio et al., 2020; Zhang et al., 2020b; Riccio and Tonella, 2020; Zhang et al., 2018; Tian et al., 2018; Pei et al., 2017; Sun et al., 2018). However, the objectives of these techniques are often quite different. Some techniques focus on finding adversarial examples (Guo et al., 2018; Zhang et al., 2020a; Kurakin et al., 2018; Croce and Hein, 2020), while other solutions aim to achieve high failure exposure and/or high values of DL-specific adequacy metrics, such as neuron (Ma et al., 2018) or surprise coverage (Kim et al., 2019), or explore the decision boundaries of the DL system (Riccio and Tonella, 2020; Kang et al., 2020).
@@
In particular, DL boundary testing targets regions of the input space where small input variations can lead to misbehaviors. Boundary inputs are crucial for evaluating the DL system’s reliability, as they often expose how it handles edge cases, transitions between operational domains, and critical decision-making regions. In traditional software systems, boundary testing is typically targeted. For example, consider a Java method sum(x, y) that adds two integers, where each parameter ranges from $-2^{32}$ to $2^{31}$. A boundary testing strategy for this method would include inputs such as the minimum allowed value ($-2^{32}$), its immediate successor ($-2^{32}+1$), an arbitrary in-range value (e.g., 100), the maximum allowed value ($2^{31}$), and its immediate predecessor ($12^{31}-$ 1). Since there are two parameters, this targeted approach yields only $255^{2}=$ 25 combinations, covering edge behaviors that are most likely to reveal bugs. In contrast, boundary testing for DL systems is challenging due to high-dimensional, unconstrained input spaces (e.g., images) and unclear input space partitions. As such, existing solutions such as DeepJanus (Riccio and Tonella, 2020) and Sinvad (Kang et al., 2024) rely on untargeted boundary testing strategies. These are commonly driven by evolutionary algorithms that generate diverse inputs without any explicit consideration of specific source or target classes. While these methods can uncover unexpected behaviors, they tend to be inefficient and unfocused, as they treat the entire input space uniformly rather than concentrating on regions near critical decision boundaries. However, DL models inherently learn decision boundaries between classes. For instance, in a digit classification task, given an image of class 5, the DL model may assign high probabilities to both the classes 5 and 6, reflecting the probabilistic nature of the model’s output rather than a definitive classification (Weiss et al., 2023). This suggests the model is uncertain between these two classes, making inputs from class 6 promising candidates for generating boundary cases. Thus, it is potentially more effective to focus testing on inputs near the classifier’s decision boundary between classes that share some features like 5 and 6, rather than sampling randomly across unrelated classes. Despite this potential, targeted boundary testing in DL systems remains largely unexplored.
@@
While researchers have explored various approaches, existing solutions have key limitations that hinder their effectiveness in boundary testing of complex DL systems. An example is DeepJanus (Riccio and Tonella, 2020), an input generation technique that relies on an abstract representation of the input domain (i.e., a model) to generate test cases. However, such domain models are typically unavailable for complex, feature-rich datasets such as ImageNet. Although recent advances in generative AI have addressed the lack of explicit input models, current techniques for generating inputs in the latent space of DL models (Kang et al., 2020; Dola et al., 2024; Dunn et al., 2020, 2021) either do not target boundary inputs, or they offer limited control over the generation process due to the use of a single, entangled latent vector perturbed by random noise (Kang et al., 2024), thereby severely constraining the ability to navigate the latent space.
@@
Our technique, implemented in a tool called Mimicry (rep, 2025), uses style-specific interpolation operations to find boundary inputs. Mimicry uses a conditional StyleGAN (Karras et al., 2019b) model trained to learn the class-wise visual characteristics of a given image dataset across all its inputs. StyleGAN maps latent vector inputs to an intermediate latent vector, which controls the image style at various granularity levels in the generative process. The main idea of Mimicry involves the systematic mutation of the pre-defined set of latent vectors between source and target inputs using scale-specific interpolation and assessing the impact of these modifications in the image space. Moreover, Mimicry facilitates the targeted generation of boundary inputs by leveraging model confidence scores. Given a source input, Mimicry identifies the boundary target as the class with the second-highest predicted confidence from the DL system. It then establishes a closed feedback loop between the DL model under test and the StyleGAN network to guide input synthesis. Specifically, Mimicry employs StyleGAN to generate representative samples of the target class and manipulates the latent representation of the source input to incorporate features of these target samples, thereby adjusting its visual characteristics toward the decision boundary.
We have evaluated the effectiveness of Mimicry on five popular image classification datasets with increasing complexity (MNIST (LeCun et al., 1998), FashionMNIST (Xiao et al., 2017), SVHN (Netzer et al., 2011), CIFAR-10 (Krizhevsky et al., 2009), ImageNet (Deng et al., 2009)) to assess its robustness across a diverse range of visual patterns and challenges, using self and pre-trained WideResNet (Zagoruyko, 2016) as DL systems under test. Additionally, we compare the effectiveness of Mimicry against the model-based DeepJanus (Riccio and Tonella, 2020) and the generative-based Sinvad (Kang et al., 2024). Our experiments demonstrate that Mimicry consistently identifies inputs close to the decision boundary while maintaining a high validity rate and label preservation rate, as evaluated by human assessors. Moreover, Mimicry surpasses both DeepJanus and Sinvad in both quantitaive and qualitative metrics, especially when increasing data complexity. Our paper makes the following contributions:
@@
To the best of our knowledge, Mimicry is the first targeted boundary testing technique for DL systems. Our approach is implemented in the publicly available tool Mimicry (rep, 2025) and is based on a disentangled latent space representation that ensures high controllability.
@@ 2.1. Testing Objectives for DL Systems
Testing methodologies to highlight behavior in DL models can have vastly different objectives. The distinctions are often unclear in the related literature. Therefore, we define key terms and specify the experimental domain. We illustrate these differences using a classifier manifold $M$, which encodes all possible classifier decisions, mapping from a high dimensional input space to the lower dimensionality space $M$. Within this manifold, distinct regions $M_{\bullet}$ (sub-manifolds) exist, corresponding to each classifiable class, respectively. In Fig. 1 the main sub-manifold for class $X$, $M_{X}$ is in focus, showing its “boundaries” to other regions on $M$ and internal boundaries which symbolize adversarial regions (Tabacof and Valle, 2016).
@@
In this work, we explore boundary testing, a subset of functional testing, which targets the generalizability aspect of the SUT by generating functionally new inputs. Some approaches such as DeepXplore (Pei et al., 2017), DLFuzz (Guo et al., 2018), and DeepTest (Tian et al., 2018) involve raw input manipulation techniques that modify/corrupt the original inputs (e.g., pixels). These techniques do not generate new functional inputs as they produce changes in the original inputs and are therefore suitable to test the deficiencies in robustness of the DL system (Riccio and Tonella, 2023; Maryam et al., 2025), such as the discovery of adversarial regions. In contrast, the generation of functional tests focuses on creating new inputs that deviate from the original training distribution. These inputs target the long-tail problem of DL testing (Zhang et al., 2024), testing the DNN’s ability to generalize to novel, unseen scenarios. Instances of functional test generators are the model-based approaches like DeepJanus (Riccio and Tonella, 2020), DeepHyperion (Zohdinasab et al., 2021) and DeepMetis (Riccio et al., 2021) or latent space manipulation techniques like SINVAD (Kang et al., 2020, 2024), CIT4DNN (Dola et al., 2024), and RBT4DNN (Mozumder et al., 2025).
@@ 2.2. Boundary Testing for DL Systems
Boundary testing identifies input samples near decision boundaries, where the classifier assigns equal, or near-equal, probabilities to multiple classes (Riccio and Tonella, 2020; Kang et al., 2020). The decision boundary of a classifer can be inferred from the predicted logits, where the theoretical boundary would be a perfect equilibrium in confidences between $n$ classes ($n>1$). In addition, boundary testing can be either targeted or untargeted. The goal of targeted boundary testing is to converge to the boundary between the origin class and a specified target class (e.g., $M_{Z}$ in 1(a)), while the goal of the untargeted case assumes no predetermined target class (1(b)).
While existing techniques such as DeepJanus (Riccio and Tonella, 2020) and Sinvad (Kang et al., 2020) focus on untargeted boundary testing, in our work, we focus on targeted boundary testing, with the goal of automatically retrieving inputs that are ambiguous in prediction, without restrictions on the input differences.
@@ 2.3. Style-Based Generative Adversarial Networks
Generative Adversarial Networks (GANs) are DL models designed to learn the statistical distribution of a training dataset, allowing the synthesis of new samples that are representative of the learned distribution (Goodfellow et al., 2020). GANs involve jointly training a pair of networks that compete with each other. This approach is based on game theory and is implemented using two neural networks. A first neural network, called the generator, aims to produce realistic images, while a second neural network, called the discriminator, acts as an expert that receives both fake and real (authentic) images and aims to distinguish between them. In this way, the generator improves its ability to produce realistic images to fool the discriminator, which can be leveraged for test generation (Dola et al., 2023; Dunn et al., 2021).
StyleGAN (Karras et al., 2019b, 2020b, 2021; Sauer et al., 2022) extends the GAN architecture to introduce new methods for controlling the image synthesis process. Unlike traditional GANs, StyleGAN enables style control at multiple levels within the network. The proposed changes to the generator model involve the use of a mapping network to map points in the initial latent space to an intermediate latent space. This intermediate latent space controls the strength of the image features at various scales in the generator model, inspired by the style transfer literature (Huang and Belongie, 2017). This architectural change, combined with the noise injected directly into the network, enables the automatic, unsupervised separation of high-level attributes from stochastic variations in the generated images, which we exploit for boundary testing.
@@ 3. Methodology
Mimicry is a black-box approach [^1] A recent survey classifies methods that need access to both the training and test datasets of a learned component as data-box methods. However, to avoid confusion, we refer to these as black-box methods, since they do not utilize any internal information from the model itself (Riccio et al., 2020). that leverages StyleGANs to generate boundary inputs through targeted optimization. Concretely, Mimicry uses four separate components (Fig. 2):
@@
Mimicry’s targeted nature lies in its treatment of boundary discovery: instead of focusing solely on maximizing misclassification (Riccio and Tonella, 2020; Kang et al., 2020, 2024), it identifies the second most probable class as the target. This initial bias toward a specific class allows Mimicry to exploit the proximity to a specific decision boundary. The target seed $w_{g}$, together with the original latent vector, is then manipulated according to a strategy $\kappa$, which is optimized by the optimizer using defined objectives $\omega_{\bullet}$. In the following sections we will describe each component of Mimicry in more detail.
@@ 3.1. System under Test
We denote the classifier as $\phi_{P}$, a trainable map from image input to the set of possible classes $C$. Specifically, the output is a vector of class probabilites $\mathbb{R}^{i}\xrightarrow{\phi_{P}}\mathbb{R}^{||C||}$.
Here, the superscript $i$ denotes the shape of the input $i$. We can think of the classification operation as positioning our input on the classifier manifold, with the location being the predicted class confidences. On this classifier manifold, Mimicry aims to finding boundaries between regions of different classes. The boundaries are regions where the classifier’s confidence $\hat{y}\in\mathbb{R}^{||C||}$ is equidistant between two or more classes in the set $C_{t}$ (missing) 1. Note here that $\sum\hat{y}=1$.
(1) $\forall c\in C_{t},\quad\hat{y}_{c}=\frac{1}{||C_{t}||}.$
@@ 3.2. Manipulator
To find boundary cases, Mimicry manipulates latent vectors from a conditional StyleGAN model. StyleGANs were specifically chosen because their disentangled latent space offers greater control over manipulations. Unlike traditional GAN architectures, where a single (noise) latent vector is used to generate outputs  (Goodfellow et al., 2020), StyleGAN uses a latent vector that passes through an additional network called the mapping network $\phi_{m}$, which consists of multiple fully connected layers. This mapping network “disentangles” the latent space by distributing learnt image characterisitcs across the layers of an intermediate latent vector $w$. The number of layers in $w$ depends on the specific StyleGAN architecture, with more layers enabling finer control over image manipulations. In contrast, traditional GANs and VAEs can be seen as having only a single such layer in $w$, which limits the degree of control.
The StyleGAN model generates new images from class information (missing) 2 and can be denoted as a composition of a mapping network $\phi_{m}$ and a synthesis network $\phi_{g}$ (missing) 3. We denote it as $\phi_{m}(\cdot)$, where the only explicitly given input is the class information, as $z$ is sampled noise.
@@
To get from a class to a generated image in StyleGAN we sample a latent vector $z\sim\mathcal{N}$ which then is processed by the mapping network in combination with the class information to generate an intermediate latent vector $w$ (missing) 4 The latent vector $w$ is then used for manipulation, as it can be separated into multiple independent layers, depending on the StyleGANs architecture. The advantage of this conversion is that the manipulation of latent vectors $z$ may produce erratic changes in the image, as observed in previous studies (Kang et al., 2020; Riccio and Tonella, 2023; Dunn et al., 2021; Dola et al., 2024).
@@
Additionally, as noted by Karras et al. (Karras et al., 2020b), the rank of manipulations in the intermediate latent space of StyleGAN2 allows for control at different levels of granularity in the generated images. Specifically, the first three layers of the intermediate latent vector $w$ tend to control coarse features such as overall shape and perspective. The next four layers influence medium-scale attributes, including textures and finer structural details. Finally, the last layers typically affect only color schemes, making them responsible for the most fine-grained manipulations. An example of these effects can be seen in Fig. 3, where layers of the original latent vector are mixed with elements of a differing target vector (here car vs. truck in CIFAR-10). While the exact layer assignments are specific to StyleGAN2, this behavior is consistent across all StyleGAN architectures, as they share the same type of intermediate latent space $w$, albeit with varying numbers of layers (Sauer et al., 2022; Karras et al., 2019b, 2020b, 2021). The intermediate latent vectors $w$ can are subsequently synthesized into an image using the synthesis network $\phi_{g}$.

### v3

@@ 1. Introduction
The increasing dependence on Deep Learning (DL) systems for both everyday tasks and critical sectors (schutze2008introduction) makes rigorous testing for these systems a relevant topic (riccio2020testing; zhang2020machine). The concept of fault in DL systems is more complex than in traditional software (riccio2020testing). Even if the code that builds the DL network is bug-free, the trained DL model may still deviate from the expected behavior due to faults introduced during the training phase, such as the misconfiguration of learning parameters or the use of an unbalanced or non-representative training set (2020-Humbatova-ICSE). In data-intensive software systems, such as DL systems, faults often stem from the large, high-dimensional input space, which requires the generation of test data that accurately captures the complexity and diversity of the validity domain, i.e., the portion of the input space for which the system is expected to operate (riccio2020testing).
@@
Test generation techniques have been developed to induce misbehaviors in DL systems (riccio2020testing; zhang2020machine; riccio2020deepjanus; deeproad; deeptest; pei2017deepxplore; sun2018testing). However, the objectives of these techniques are often quite different. Some techniques focus on finding adversarial examples (guo2018dlfuzz; zhang2020deepsearch; kurakin2018adversarial; croce2020minimally), while other solutions aim to achieve high failure exposure and/or high values of DL-specific adequacy metrics, such as neuron (ma2018deepgauge) or surprise coverage (kim2019guiding), or explore the decision boundaries of the DL system (riccio2020deepjanus; kang2020sinvad).
@@
In particular, DL boundary testing targets regions of the input space where small input variations can lead to misbehaviors. Boundary inputs are crucial for evaluating the DL system’s reliability, as they often expose how it handles edge cases, transitions between operational domains, and critical decision-making regions. In traditional software systems, boundary testing is typically targeted. For example, consider a Java method sum(x, y) that adds two integers, where each parameter ranges from $2^{32}$ to $2^{31}$. A boundary testing strategy for this method would include inputs such as the minimum allowed value ($2^{32}$), its immediate successor ($2^{32}+1$), an arbitrary in-range value (e.g., 100), the maximum allowed value ($2^{31}$), and its immediate predecessor ($2^{31}-1$). Since there are two parameters, this targeted approach yields only $5^{2}=25$ combinations, covering edge behaviors that are most likely to reveal bugs. In contrast, boundary testing for DL systems is challenging due to high-dimensional, unconstrained input spaces (e.g., images) and unclear input space partitions. As such, existing solutions such as DeepJanus (riccio2020deepjanus) and Sinvad (sinvad-tosem) rely on untargeted boundary testing strategies. These are commonly driven by evolutionary algorithms that generate diverse inputs without any explicit consideration of specific source or target classes. While these methods can uncover unexpected behaviors, they tend to be inefficient and unfocused, as they treat the entire input space uniformly rather than concentrating on regions near critical decision boundaries. However, DL models inherently learn decision boundaries between classes. For instance, in a digit classification task, given an image of class 5, the DL model may assign high probabilities to both the classes 5 and 6, reflecting the probabilistic nature of the model’s output rather than a definitive classification (10.1007/s10664-023-10393-w). This suggests the model is uncertain between these two classes, making inputs from class 6 promising candidates for generating boundary cases. Thus, it is potentially more effective to focus testing on inputs near the classifier’s decision boundary between classes that share some features like 5 and 6, rather than sampling randomly across unrelated classes. Despite this potential, targeted boundary testing in DL systems remains largely unexplored.
@@
While researchers have explored various approaches, existing solutions have key limitations that hinder their effectiveness in boundary testing of complex DL systems. An example is DeepJanus (riccio2020deepjanus), an input generation technique that relies on an abstract representation of the input domain (i.e., a model) to generate test cases. However, such domain models are typically unavailable for complex, feature-rich datasets such as ImageNet. Although recent advances in generative AI have addressed the lack of explicit input models, current techniques for generating inputs in the latent space of DL models (kang2020sinvad; dola2024cit4dnn; DBLP:journals/corr/abs-2001-11055; dunn2021exposing) either do not target boundary inputs, or they offer limited control over the generation process due to the use of a single, entangled latent vector perturbed by random noise (sinvad-tosem), thereby severely constraining the ability to navigate the latent space.
@@
Our technique, implemented in a tool called Mimicry (replication-package), uses style-specific interpolation operations to find boundary inputs. Mimicry uses a conditional StyleGAN (karras2019style) model trained to learn the class-wise visual characteristics of a given image dataset across all its inputs. StyleGAN maps latent vector inputs to an intermediate latent vector, which controls the image style at various granularity levels in the generative process. The main idea of Mimicry involves the systematic mutation of the pre-defined set of latent vectors between source and target inputs using scale-specific interpolation and assessing the impact of these modifications in the image space. Moreover, Mimicry facilitates the targeted generation of boundary inputs by leveraging model confidence scores. Given a source input, Mimicry identifies the boundary target as the class with the second-highest predicted confidence from the DL system. It then establishes a closed feedback loop between the DL model under test and the StyleGAN network to guide input synthesis. Specifically, Mimicry employs StyleGAN to generate representative samples of the target class and manipulates the latent representation of the source input to incorporate features of these target samples, thereby adjusting its visual characteristics toward the decision boundary.
We have evaluated the effectiveness of Mimicry on five popular image classification datasets with increasing complexity (MNIST (mnist), FashionMNIST (fashionmnist), SVHN (svhn), CIFAR-10 (cifar), ImageNet (deng2009imagenet)) to assess its robustness across a diverse range of visual patterns and challenges, using self and pre-trained WideResNet (zagoruyko2016wide) as DL systems under test. Additionally, we compare the effectiveness of Mimicry against the model-based DeepJanus (riccio2020deepjanus) and the generative-based Sinvad (sinvad-tosem). Our experiments demonstrate that Mimicry consistently identifies inputs close to the decision boundary while maintaining a high validity rate and label preservation rate, as evaluated by human assessors. Moreover, Mimicry surpasses both DeepJanus and Sinvad in both quantitaive and qualitative metrics, especially when increasing data complexity. Our paper makes the following contributions:
@@
To the best of our knowledge, Mimicry is the first targeted boundary testing technique for DL systems. Our approach is implemented in the publicly available tool Mimicry (replication-package) and is based on a disentangled latent space representation that ensures high controllability.
@@ 2.1. Testing Objectives for DL Systems
Testing methodologies to highlight behavior in DL models can have vastly different objectives. The distinctions are often unclear in the related literature. Therefore, we define key terms and specify the experimental domain. We illustrate these differences using a classifier manifold $M$, which encodes all possible classifier decisions, mapping from a high dimensional input space to the lower dimensionality space $M$. Within this manifold, distinct regions $M_{\bullet}$ (sub-manifolds) exist, corresponding to each classifiable class, respectively. In Fig. 1 the main sub-manifold for class $X$, $M_{X}$ is in focus, showing its “boundaries” to other regions on $M$ and internal boundaries which symbolize adversarial regions (7727230).
@@
In this work, we explore boundary testing, a subset of functional testing, which targets the generalizability aspect of the SUT by generating functionally new inputs. Some approaches such as DeepXplore (pei2017deepxplore), DLFuzz (guo2018dlfuzz), and DeepTest (deeptest) involve raw input manipulation techniques that modify/corrupt the original inputs (e.g., pixels). These techniques do not generate new functional inputs as they produce changes in the original inputs and are therefore suitable to test the deficiencies in robustness of the DL system (2023-Riccio-ICSE; 2025-Maryam-ICST), such as the discovery of adversarial regions. In contrast, the generation of functional tests focuses on creating new inputs that deviate from the original training distribution. These inputs target the long-tail problem of DL testing (zhang2024systematicreviewlongtailedlearning), testing the DNN’s ability to generalize to novel, unseen scenarios. Instances of functional test generators are the model-based approaches like DeepJanus (riccio2020deepjanus), DeepHyperion (zohdinasab2021deephyperion) and DeepMetis (2021-Riccio-ASE) or latent space manipulation techniques like SINVAD (kang2020sinvad; sinvad-tosem), CIT4DNN (dola2024cit4dnn), and RBT4DNN (mozumder2025rbt4dnnrequirementsbasedtestingneural).
@@ 2.2. Boundary Testing for DL Systems
Boundary testing identifies input samples near decision boundaries, where the classifier assigns equal, or near-equal, probabilities to multiple classes (riccio2020deepjanus; kang2020sinvad). The decision boundary of a classifer can be inferred from the predicted logits, where the theoretical boundary would be a perfect equilibrium in confidences between $n$ classes ($n>1$). In addition, boundary testing can be either targeted or untargeted. The goal of targeted boundary testing is to converge to the boundary between the origin class and a specified target class (e.g., $M_{Z}$ in 1(a)), while the goal of the untargeted case assumes no predetermined target class (1(b)).
While existing techniques such as DeepJanus (riccio2020deepjanus) and Sinvad (kang2020sinvad) focus on untargeted boundary testing, in our work, we focus on targeted boundary testing, with the goal of automatically retrieving inputs that are ambiguous in prediction, without restrictions on the input differences.
@@ 2.3. Style-Based Generative Adversarial Networks
Generative Adversarial Networks (GANs) are DL models designed to learn the statistical distribution of a training dataset, allowing the synthesis of new samples that are representative of the learned distribution (goodfellow2020generative). GANs involve jointly training a pair of networks that compete with each other. This approach is based on game theory and is implemented using two neural networks. A first neural network, called the generator, aims to produce realistic images, while a second neural network, called the discriminator, acts as an expert that receives both fake and real (authentic) images and aims to distinguish between them. In this way, the generator improves its ability to produce realistic images to fool the discriminator, which can be leveraged for test generation (dola2023input26; dunn2021exposing).
StyleGAN (karras2019style; karras2020analyzing; karras2021alias; sauer2022styleganXL) extends the GAN architecture to introduce new methods for controlling the image synthesis process. Unlike traditional GANs, StyleGAN enables style control at multiple levels within the network. The proposed changes to the generator model involve the use of a mapping network to map points in the initial latent space to an intermediate latent space. This intermediate latent space controls the strength of the image features at various scales in the generator model, inspired by the style transfer literature (Huang_2017_ICCV). This architectural change, combined with the noise injected directly into the network, enables the automatic, unsupervised separation of high-level attributes from stochastic variations in the generated images, which we exploit for boundary testing.
@@ 3. Methodology
Mimicry is a black-box approach [^1] A recent survey classifies methods that need access to both the training and test datasets of a learned component as data-box methods. However, to avoid confusion, we refer to these as black-box methods, since they do not utilize any internal information from the model itself (riccio2020testing). that leverages StyleGANs to generate boundary inputs through targeted optimization. Concretely, Mimicry uses four separate components (Fig. 2):
@@
Mimicry’s targeted nature lies in its treatment of boundary discovery: instead of focusing solely on maximizing misclassification (riccio2020deepjanus; kang2020sinvad; sinvad-tosem), it identifies the second most probable class as the target. This initial bias toward a specific class allows Mimicry to exploit the proximity to a specific decision boundary. The target seed $w_{g}$, together with the original latent vector, is then manipulated according to a strategy $\kappa$, which is optimized by the optimizer using defined objectives $\omega_{\bullet}$. In the following sections we will describe each component of Mimicry in more detail.
@@ 3.1. System under Test
[…]
@@ 3.2. Manipulator
To find boundary cases, Mimicry manipulates latent vectors from a conditional StyleGAN model. StyleGANs were specifically chosen because their disentangled latent space offers greater control over manipulations. Unlike traditional GAN architectures, where a single (noise) latent vector is used to generate outputs  (goodfellow2020generative), StyleGAN uses a latent vector that passes through an additional network called the mapping network $\phi_{m}$, which consists of multiple fully connected layers. This mapping network “disentangles” the latent space by distributing learnt image characterisitcs across the layers of an intermediate latent vector $w$. The number of layers in $w$ depends on the specific StyleGAN architecture, with more layers enabling finer control over image manipulations. In contrast, traditional GANs and VAEs can be seen as having only a single such layer in $w$, which limits the degree of control.
The StyleGAN model generates new images from class information () 2 and can be denoted as a composition of a mapping network $\phi_{m}$ and a synthesis network $\phi_{g}$ () 3. We denote it as $\phi_{m}(\cdot)$, where the only explicitly given input is the class information, as $z$ is sampled noise.
@@
To get from a class to a generated image in StyleGAN we sample a latent vector $z\sim\mathcal{N}$ which then is processed by the mapping network in combination with the class information to generate an intermediate latent vector $w$ () 4 The latent vector $w$ is then used for manipulation, as it can be separated into multiple independent layers, depending on the StyleGANs architecture. The advantage of this conversion is that the manipulation of latent vectors $z$ may produce erratic changes in the image, as observed in previous studies (kang2020sinvad; 2023-Riccio-ICSE; dunn2021exposing; dola2024cit4dnn).
@@
Additionally, as noted by Karras et al. (karras2020analyzing), the rank of manipulations in the intermediate latent space of StyleGAN2 allows for control at different levels of granularity in the generated images. Specifically, the first three layers of the intermediate latent vector $w$ tend to control coarse features such as overall shape and perspective. The next four layers influence medium-scale attributes, including textures and finer structural details. Finally, the last layers typically affect only color schemes, making them responsible for the most fine-grained manipulations. An example of these effects can be seen in Fig. 3, where layers of the original latent vector are mixed with elements of a differing target vector (here car vs. truck in CIFAR-10). While the exact layer assignments are specific to StyleGAN2, this behavior is consistent across all StyleGAN architectures, as they share the same type of intermediate latent space $w$, albeit with varying numbers of layers (sauer2022styleganXL; karras2019style; karras2020analyzing; karras2021alias). The intermediate latent vectors $w$ can are subsequently synthesized into an image using the synthesis network $\phi_{g}$.